EXPLORING DYNAMICAL EFFECTS OF BEND EXCITATION ON THE Cl + CH_{4} REACTION

\underline{HANS A. BECHTEL}^{a}, JON P. CAMDEN, DAVIDA J. ANKENY BROWN, MARION R. MARTIN, RICHARD N. ZARE, Department of Chemistry, Stanford University, Stanford, California 94305-5080; and KONSTANTIN VODOPYANOV, Ginzton Laboratory, Stanford University, Stanford, California 94305-4088.

Using a novel IR source, we have directly pumped the CH_{4}(ν_{4}=1) and CH_{3}D(ν_{3}=1) bending vibrations and detected the nascent CH_{3} products from the Cl + CH_{4} reaction with 2+1 REMPI in a time-of-flight mass spectrometer. Contrary to intuitive expectations, we find that these low-frequency bending vibrations enhance the cross section for reaction with Cl atoms by a factor of two or more. This enhancement is constant over the collision energy range 0.12 eV to 0.26 eV. Moreover, the nonlocalized motion in the methane reagent is transferred almost exclusively to translational motion of the escaping products, which is in sharp contrast to theoretical calculations that predict formation of methyl radicals predominantly excited into the umbrella bending mode. We also observe both H-atom and D-atom abstraction products from the Cl + CH_{3}D(ν_{3}=1) reaction, indicating that the reaction is not bond selective.

aPresent address: Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139